OpenStudy (anonymous):

Solve the following using the quadratic formula. Round to the nearest tenth, if necessary. x2 + 5x - 2 = 0 {-2, 5} {-5.4, 0.4} {-1, 0.5} No real solution

OpenStudy (mertsj):

Do you know the quadratic formula?

OpenStudy (mertsj):

Look in you book. It's in there.

OpenStudy (mertsj):

Did you find it?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So in your equation, a = 1, b = 5, and c = -2 Just plug in those numbers in place of the letters.

OpenStudy (mertsj):

\[x=\frac{-5\pm \sqrt{5^2-4(1)(-2)}}{2(1)}\]

OpenStudy (mertsj):

Does yours look like that so far?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[x=\frac{-5\pm \sqrt{33}}{2}\]

OpenStudy (mertsj):

Oh gees. They gave the answers in decimal form so now we need a calculator. Do you have one handy?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[\frac{-5+\sqrt{33}}{2}=.37\approx .4\] \[\frac{-5-\sqrt{33}}{2}=-5.37\approx -5.4\]

OpenStudy (mertsj):

Is that what you got?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Good job!!

OpenStudy (anonymous):

Thanks:)

OpenStudy (mertsj):

you're welcome